# AccidentBlip2: Accident Detection With Multi-View MotionBlip2

### v1

@@ IV-D Evaluation
$\checkmark$ $\checkmark$ -- $\checkmark$ 66.7
@@
Our proposed model outperform baseline models in both single-vehicle and multi-vehicle configurations. Compared with these video large language models, shown in table II, our AccidentBlip2 achieves 66.5% with merely single car input. In comparison with V2XFormer, proposed with three different V2X-agent fusion setting, a significant increment can be observed in our model, with about 3% improvement in accuracy.
TABLE II: Performance Comparison Amongst Different Video Varge Language Models.
Model Data Sources Accuracy
Video-LLaMA single vehicle 51.34
multiple vehicles 33.33
Video-Vicuna single vehicles 36.43
multiple vehicles 31.87
Video-LLaVA single vehicle 49.7
multiple vehicles 47.7
V2XFormer single vehicle 43.8
4 vehicles + infra 69.5
AccidentBlip2 (ours) single vehicle 66.5
multiple vehicles 73.1
Our approach is designed to address the challenge of gradient shifts during training, which can lead to gradient explosion in the initial epochs. To mitigate this issue, we employ a warmup scheduler that makes slight adjustments to the learning rate within a minimal range. This helps stabilize the training process. Our multi-vehicle paradigm operates on four separate GPUs, allowing us to gather query results from each motion Qformer. Additionally, we conducted experiments to explore the impact of the initial weight of the Qformer, comparing pre-trained weights from Blip2 with default initialization. The results revealed a decrease in the overall performance of our language model when evaluated on the DeepAccident dataset. To overcome this, we utilize pre-trained weights from Blip2 to initialize AccidentBlip2, improving the model’s effectiveness in accident detection tasks.

### v4

@@ IV-C Main Result
$\checkmark$ $\checkmark$ -- $\checkmark$ 70.4
@@
Analysis on multi-vehicle paradigm. Our approach is designed to address the challenge of gradient shifts during training, which can lead to gradient explosion in the initial epochs. To mitigate this issue, we employ a warmup scheduler that makes slight adjustments to the learning rate within a minimal range. This helps stabilize the training process. Our multi-vehicle paradigm operates on four separate GPUs, allowing us to gather query results from each motion Qformer. Additionally, we conducted experiments to explore the impact of the initial weight of the Qformer, comparing pre-trained weights from Blip2 with default initialization. The results revealed a decrease in the overall performance of our language model when evaluated on the DeepAccident dataset. To overcome this, we utilize pre-trained weights from Blip2 to initialize AccidentBlip2, improving the model’s effectiveness in accident detection tasks.